# Flashlights: Prospects for constraining the Initial Mass Function around cosmic noon with caustic-crossing events

### v1

@@
Abstract
The Flashlights program with the Hubble Space Telescope imaged the six Hubble Frontier Fields galaxy clusters in two epochs and detected twenty transients. These are primarily expected to be caustic-crossing events (CCEs) where bright stars in distant lensed galaxies, typically at redshift $z\approx 1$–3, get temporarily magnified close to cluster caustics. Since CCEs are generally biased toward more massive and luminous stars, they offer a unique route for probing the high end of the stellar mass function. We take advantage of the Flashlights event statistics to place preliminary constraints on the stellar initial mass function (IMF) around cosmic noon. The photometry (along with spectral information) of lensed arcs is used to infer their various stellar properties, and stellar synthesis models are used to evolve a recent stellar population in them. We estimate the microlens surface density near each arc and, together with existing lens models and simple formalism for CCEs, calculate the expected rate for a given IMF. We find that, on average, a Salpeter-like IMF ($\alpha=2.35$) underpredicts the number of observed CCEs by a factor of ${\sim}0.7$, and a top-heavy IMF ($\alpha=1.00$) overpredicts by a factor of ${\sim}1.7$, suggesting that the average IMF slope may lie somewhere in between. However, given the large uncertainties associated with estimating the stellar populations, these results are strongly model-dependent. Nevertheless, we introduce a useful framework for constraining the IMF using CCEs. Observations with JWST are already yielding many more CCEs and will soon enable more stringent constraints on the IMF at a range of redshifts.
@@ Abstract
sf
thanks: ashishmeena766@gmail.com
@@ 1 Introduction
Understanding the formation and evolution of stars and galaxies across cosmic history is one of the main interests in astronomy (e.g., Conselice, 2014; Madau & Dickinson, 2014). Giant, supergiant and hypergiant stars (which we shall call throughout “supergiants" for simplicity) dominate the light coming from young and star-forming galaxies. Such stars only live for several million years (Myr) and end their life through supernovae, further affecting the properties of their host galaxies (e.g., Eldridge & Stanway, 2012, 2022). In general, we can only observe and study individual supergiant stars directly within our own Galaxy or its vicinity (e.g., Evans et al., 2007; Castro et al., 2018), while for the more distant universe (e.g., redshift $z\gtrsim 1$), we rely on integrated light or comparison with synthesized stellar population models to constrain their properties. Our understanding of the formation and evolution of supergiant stars and their effects on the integrated light of the host galaxy is thus still far from complete (e.g., Conroy, 2013; Eldridge et al., 2017). One of the key missing pieces of the puzzle is their abundance; for example, various studies suggest that there may be significantly more blue supergiants than predicted by classical stellar evolution theory (sometimes referred to as the “blue supergiant problem"; e.g., Bellinger et al. 2024).
@@
Given that supergiants are short-lived and thus probe recent phases of star formation, their abundance in different galaxies through cosmic history can also help assess the universality – or evolution – of the stellar initial mass function (IMF), especially on its high-mass end, which would in turn have strong implications for galaxy evolution models. This may have become even more timely now, in light of recent results from JWST. For example, some JWST results suggest an overabundance of ultraviolet (UV)-bright galaxies at cosmic dawn (e.g., Chemerynska et al., 2024; McLeod et al., 2024), and sufficient contribution by faint galaxies to the reionization of the Universe (Atek et al., 2024, c.f. Naidu et al. 2020), where most of the UV radiation in them is believed to come from hot, massive stars. Some early galaxies have been found to show hints of early mass build-up (Labbé et al., 2023; Boyett et al., 2024), which may be explained by very efficient star-formation channels (e.g., Dekel et al., 2023) or top-heavier IMFs (e.g., Steinhardt et al., 2022).
@@
On average, galaxies decrease in size and brightness with redshift (e.g., Bouwens et al., 2004; Holwerda et al., 2015; Kawamata et al., 2015; Mowla et al., 2019; Ormerod et al., 2024; Donnan et al., 2023), but thanks to gravitational lensing magnification, some background galaxies, especially behind galaxy clusters, can get magnified substantially, revealing small-scale details within them (e.g., Whitaker et al., 2014; Rigby et al., 2018; Adamo et al., 2024; Fujimoto et al., 2024). A few years ago, lensing opened a route to observing individual stars at cosmological distances (Kelly et al., 2018). When stars in strongly magnified background galaxies that sit atop (or near) the lensing critical curves of a galaxy cluster lens, where the magnification gets extremely high, sweep across the net of microcaustics created by the microlenses in the intracluster medium (ICM), they get sufficiently magnified to be observed for short periods of time. Such events are known as caustic transients or caustic-crossing events (CCEs; Miralda-Escude, 1991; Venumadhav et al., 2017; Diego et al., 2018; Oguri et al., 2018). Roughly 80–90 (confirmed and candidates) of these CCEs have by now been detected with the Hubble Space Telescope (HST) and JWST, at various redshifts all the way to the first billion years in the Universe (e.g., Rodney et al., 2018; Chen et al., 2019; Kaurov et al., 2019; Diego et al., 2022; Kelly et al., 2022; Diego et al., 2023; Meena et al., 2023a, b; Yan et al., 2023; Diego et al., 2024a; Fudamoto et al., 2024) with numbers continuously increasing. In addition to observing these CCE events in HST and JWST images, spectra have been obtained of some of these lensed star candidates (Diego et al., 2022; Furtak et al., 2024; Pascale & Dai, 2024; Choe et al., 2024).
@@
The peak magnification a star obtains when crossing a microcaustic depends on various factors such as the radius of the star, underlying (macro-) magnification, and the microlens density along with its mass function (e.g., Venumadhav et al., 2017; Oguri et al., 2018). High-cadence observations around these peaks can even reveal intrinsic properties of the background individual stars (e.g., Han & Dai, 2024). CCEs are expected mostly close to the macrocritical curves (usually $\lesssim 0.3^{\prime\prime}$ but can go up to $\sim 1^{\prime\prime}$ for low-redshift lensed arcs such as the Dragon arc in Abell 370 galaxy cluster; Fudamoto et al. 2024), where the underlying macromagnification is sufficiently high, on the order of $\gtrsim 10^{2}$. However, very close to the macrocritical curve, there forms a so-called corrugated network of microcaustics such that the peak magnification remains on average constant, and lensed star candidates lying within it can also be observed as (semi-)persistent sources that remain observable for longer periods of time (e.g., Welch et al., 2022a, b; Meena et al., 2023b). Farther away from the macrocritical curve, the macromagnification decreases (as $\mu_{\rm m}\propto 1/d$, where $d$ is the distance from the macrocritical curve) along with the probability of observing CCEs. Nevertheless, the smaller the magnification, the larger the area, and the stronger also is the bias toward observing more-luminous stars as they require less magnification to be observed. Hence, CCEs are an excellent route for tracing and for constraining the number density of this intrinsically luminous population of stars in the distant Universe and, thus, for constraining their IMF (Windhorst et al., 2018). In addition, since the event rate depends on the mass function and surface density of microlenses, CCEs also constitute a useful means for understanding the composition of dark matter in the intracluster medium (ICM; e.g., Dai et al., 2018; Diego et al., 2018; Oguri et al., 2018; Diego et al., 2024b; Broadhurst et al., 2025; Müller & Miralda-Escudé, 2025).
@@
In our current work, we use the event statistics of CCEs detected in Flashlights – an HST program designed to detect CCEs and other transients (Kelly et al., 2022) – to put preliminary constraints on the IMF in lensed galaxies around cosmic noon ($z\approx 1$–3). Given that this is a relatively new field and the number of CCEs detected by Flashlights is relatively small, we study here the effect of two different IMFs (parameterised as  $dn/dm\propto m^{-\alpha}$), top-heavy ($\alpha=1.0$) and Salpeter ($\alpha=2.35$; Salpeter 1955), on the expected number of CCEs. We focus on seventeen different lensed arcs, seven of which lead to the twenty CCEs observed in Flashlights as reported by Kelly et al. (2022). In addition to presenting the formalism to calculate the number of CCEs, we discuss various sources of uncertainty. Future work, including more events (which also helps reduce the related uncertainties), could then use a similar framework to put more-stringent constraints on the IMF.
@@ 2 Data and Observations
All of the HFF clusters were earlier also observed under the Frontier Fields Survey program in three HST-ACS/WFC (F435W, F606W, and F814W) and four HST-WFC3/IR (F105W, F125W, F140W, and F160W) filters, complementing previous observations and reaching a $5\sigma$ limiting magnitude of ${\sim}29$ AB (Lotz et al., 2017). We use these observations to measure the photometry and derive relevant stellar properties for each arc, and to estimate the microlens density in the vicinity of the arcs. For these purposes we use the Bayesian Analysis of Galaxies for Physical Inference and Parameter EStimation (bagpipes; Carnall et al., 2018) and spisea (Hosek et al., 2020) codes. More details are given in the section below.
Figure 1: Lensed arcs studied in our current work. In each panel, the name of the galaxy cluster, lensed arc system number (same as shown in Table 1), and arc redshift are given in the upper-left part. The red curves show the critical curve at the arc redshift corresponding to the lens model (mentioned in the upper-right part of the panel) used in the current work for various estimates for each arc. The green rectangles mark apertures used to estimate the ICL density near each arc. The color images are made using HFF data with R=F105W+F125W+140W+160W; G=F606W+F814W+F105W; B=F435W+F606W.
3 Methodology to Estimate CCE Rate
@@
3.1 The Sample of Lensed Arcs
Lensed arcs used in the current work are listed in Table 1, and the corresponding color stamps are shown in Figure 1. We study a total of seventeen cluster caustic-crossing arcs in Flashlights. The twenty transients detected in the Flashlights program during the two visits are only found in seven of these arcs, whereas the remaining ten did not yield any transients. We refer readers to Kelly et al. (2022) for a view of the Flashlights difference imaging showing the transients in the various lensed arcs.
While this does not affect our analysis or results, we note that this is not a complete list of lensed arcs around cosmic noon in the HFF clusters. For example, there are some other lensed arcs that lie very close to one of the cluster galaxies such that they are heavily contaminated (complicating the analysis and likely also reducing the chances for observing CCEs in them), or pairs of lensed galaxy images that do not necessarily form arcs but can have certain parts with magnification factors $>10$ which may be sufficient to give rise to observable CCEs from very bright stars. Although their contribution is not expected to be major, a more exhaustive analysis should also consider such systems and is left for future work.
@@ 3.2 Microlens Density
Before estimating the stellar population in the source galaxies, one important ingredient for calculating the rate of CCEs is the microlens density in the ICM close to these arcs. We place multiple rectangular apertures near each arc (see Figure 1) to photometrically measure the intracluster light (ICL) in the seven HFF filters. We use bagpipes to fit a spectral energy distribution (SED) to the photometry and estimate the ICL surface density near the lensed arcs. bagpipes uses simple stellar population synthesis (SSPS) models with the Kroupa IMF (Kroupa & Boily, 2002) along with an exponentially declining star-formation history (SFH) to do this. The stellar microlens density we measure near each lensed arc is shown in column (5) of Table 1, and we use the best-fit values for various estimations.

### v2

@@
11institutetext: Department of Physics, Ben-Gurion University of the Negev, PO Box 653, Beer-Sheva 84105, Israel 22institutetext: Department of Physics, The University of Hong Kong, Pokfulam Road, Hong Kong 33institutetext: School of Physics and Astronomy, University of Minnesota, 116 Church Street SE, Minneapolis, MN 55455, USA 44institutetext: Donostia International Physics Center, DIPC, Basque Country, San Sebastián, 20018, Spain 55institutetext: Department of Physics, University of Basque Country UPV/EHU, Bilbao, Spain 66institutetext: Ikerbasque, Basque Foundation for Science, Bilbao, Spain 77institutetext: Department of Physics, Oklahoma State University, 145 Physical Sciences Bldg, Stillwater, OK 74078, USA 88institutetext: IFCA, Instituto de F´ısica de Cantabria (UC-CSIC), Av. de Los Castros s/n, 39005 Santander, Spain 99institutetext: Department of Astronomy, University of California, Berkeley, CA 94720-3411, USA
Flashlights: Prospects for constraining the initial mass function around cosmic noon with caustic-crossing events
@@
The Flashlights program with the Hubble Space Telescope imaged the six Hubble Frontier Fields galaxy clusters in two epochs and detected twenty transients. These are primarily expected to be caustic-crossing events (CCEs) where bright stars in distant lensed galaxies, typically at redshift $z\approx 1$–3, get temporarily magnified close to cluster caustics. Since CCEs are generally biased toward more massive and luminous stars, they offer a unique route for probing the high end of the stellar mass function. We take advantage of the Flashlights event statistics to place preliminary constraints on the stellar initial mass function (IMF) around cosmic noon. The photometry (along with spectral information) of lensed arcs is used to infer their various stellar properties, and stellar synthesis models are used to evolve a recent stellar population in them. We estimate the microlens surface density near each arc and, together with existing lens models and simple formalism for CCEs, calculate the expected rate for a given IMF. We find that, on average, a Salpeter-like IMF ($\alpha=2.35$) underpredicts the number of observed CCEs by a factor of ${\sim}0.7$, and a top-heavy IMF ($\alpha=1.00$) overpredicts by a factor of ${\sim}1.7$, suggesting that the average IMF slope may lie somewhere in between. However, given the large uncertainties associated with estimating the stellar populations, these results are strongly model-dependent. Nevertheless, we introduce a useful framework for constraining the IMF using CCEs. Observations with James Webb Space Telescope are already yielding many more CCEs and will soon enable more stringent constraints on the IMF at a range of redshifts.
Key Words.:
Gravitational lensing: strong – Gravitational lensing: micro – Stars: mass function
@@ 1 Introduction
Understanding the formation and evolution of stars and galaxies across cosmic history is one of the main interests in astronomy (e.g., Conselice, 2014; Madau & Dickinson, 2014). Giant, supergiant, and hypergiant stars (which we shall call throughout “supergiants” for simplicity) dominate the light coming from young and star-forming galaxies. Such stars only live for several million years (Myr) and end their life through supernovae, further affecting the properties of their host galaxies (e.g., Eldridge & Stanway, 2012, 2022). In general, we can only observe and study individual supergiant stars directly within our own Galaxy or its vicinity (e.g., Evans et al., 2007; Castro et al., 2018), while for the more distant universe (e.g., redshift $z\gtrsim 1$), we rely on integrated light or comparison with synthesized stellar population models to constrain their properties. Our understanding of the formation and evolution of supergiant stars and their effects on the integrated light of the host galaxy is thus still far from complete (e.g., Conroy, 2013; Eldridge et al., 2017). One of the key missing pieces of the puzzle is their abundance; for example, various studies suggest that there may be significantly more blue supergiants than predicted by classical stellar evolution theory (sometimes referred to as the “blue supergiant problem”; e.g., Bellinger et al. 2024).
@@
Given that supergiants are short-lived and thus probe recent phases of star formation, their abundance in different galaxies through cosmic history can also help assess the universality – or evolution – of the stellar initial mass function (IMF), especially on its high-mass end, which would in turn have strong implications for galaxy evolution models. This may have become even more timely now, in light of recent results from James Webb Space Telescope (JWST). For example, some JWST results suggest an overabundance of ultraviolet (UV)-bright galaxies at cosmic dawn (e.g., Chemerynska et al., 2024; McLeod et al., 2024), and sufficient contribution by faint galaxies to the reionization of the Universe (Atek et al., 2024, c.f. Naidu et al. 2020), where most of the UV radiation in them is believed to come from hot, massive stars. Some early galaxies have been found to show hints of early mass build-up (Labbé et al., 2023; Boyett et al., 2024), which may be explained by very efficient star-formation channels (e.g., Dekel et al., 2023) or top-heavier IMFs (e.g., Steinhardt et al., 2022).
@@
On average, galaxies decrease in size and brightness with redshift (e.g., Bouwens et al., 2004; Holwerda et al., 2015; Kawamata et al., 2015; Mowla et al., 2019; Ormerod et al., 2024; Donnan et al., 2023), but thanks to gravitational lensing magnification, some background galaxies, especially behind galaxy clusters, can get magnified substantially, revealing small-scale details within them (e.g., Whitaker et al., 2014; Rigby et al., 2018; Adamo et al., 2024; Fujimoto et al., 2024). A few years ago, lensing opened a route to observing individual stars at cosmological distances (Kelly et al., 2018). When stars in strongly magnified background galaxies that sit atop (or near) the lensing critical curves of a galaxy cluster lens, where the magnification gets extremely high, sweep across the net of microcaustics created by the microlenses in the intracluster medium (ICM), they get sufficiently magnified to be observed for short periods of time. Such events are known as caustic transients or caustic-crossing events (CCEs; Miralda-Escude, 1991; Venumadhav et al., 2017; Diego et al., 2018; Oguri et al., 2018). Roughly 80–90 (confirmed and candidates) of these CCEs have by now been detected with the Hubble Space Telescope (HST) and JWST, at various redshifts all the way to the first billion years in the Universe (e.g., Rodney et al., 2018; Chen et al., 2019; Kaurov et al., 2019; Diego et al., 2022; Kelly et al., 2022; Diego et al., 2023; Meena et al., 2023a, b; Yan et al., 2023; Diego et al., 2024b; Fudamoto et al., 2025) with numbers continuously increasing. In addition to observing these CCE events in HST and JWST images, spectra have been obtained of some of these lensed star candidates (Diego et al., 2022; Furtak et al., 2024; Pascale & Dai, 2024; Choe et al., 2025).
@@
The peak magnification a star obtains when crossing a microcaustic depends on various factors such as the radius of the star, underlying (macro-) magnification, and the microlens density along with its mass function (e.g., Venumadhav et al., 2017; Oguri et al., 2018). High-cadence observations around these peaks can even reveal intrinsic properties of the background individual stars (e.g., Han & Dai, 2024). CCEs are expected mostly close to the macrocritical curves (usually $\lesssim 0.3^{\prime\prime}$ but can go up to $\sim 1^{\prime\prime}$ for low-redshift lensed arcs such as the Dragon arc in Abell 370 galaxy cluster; Fudamoto et al. 2025), where the underlying macromagnification is sufficiently high, on the order of $\gtrsim 10^{2}$. However, very close to the macrocritical curve, there forms a so-called corrugated network of microcaustics such that the peak magnification remains on average constant, and lensed star candidates lying within it can also be observed as (semi-) persistent sources that remain observable for longer periods of time (e.g., Welch et al., 2022a, b; Meena et al., 2023b). Farther away from the macrocritical curve, the macromagnification decreases (as $\mu_{\rm m}\propto 1/d$, where $d$ is the distance from the macrocritical curve) along with the probability of observing CCEs. Nevertheless, the smaller the magnification, the larger the area, and the stronger also is the bias toward observing more luminous stars as they require less magnification to be observed. Hence, CCEs are an excellent route for tracing and for constraining the number density of this intrinsically luminous population of stars in the distant Universe and, thus, for constraining their IMF (Windhorst et al., 2018). In addition, since the event rate depends on the mass function and surface density of microlenses, CCEs also constitute a useful means for understanding the composition of dark matter in the intracluster medium (ICM; e.g., Dai et al., 2018; Diego et al., 2018; Oguri et al., 2018; Diego et al., 2024a; Broadhurst et al., 2025; Müller & Miralda-Escudé, 2025).
@@
In our current work, we use the event statistics of CCEs detected in Flashlights – an HST program designed to detect CCEs and other transients (Kelly et al., 2022) – to put preliminary constraints on the IMF in lensed galaxies around cosmic noon ($z\approx 1$–3). Given that this is a relatively new field and the number of CCEs detected by Flashlights is relatively small, we study here the effect of two different IMFs (parameterized as  $dn/dm\propto m^{-\alpha}$), top-heavy ($\alpha=1.0$) and Salpeter ($\alpha=2.35$; Salpeter 1955), on the expected number of CCEs. We focus on seventeen different lensed arcs, seven of those lead to the twenty CCEs observed in Flashlights as reported by Kelly et al. (2022). In addition to presenting the formalism to calculate the number of CCEs, we discuss various sources of uncertainty. Future work, including more events (which also helps reduce the related uncertainties), could then use a similar framework to put more stringent constraints on the IMF.
@@ 2 Data and observations
All of the HFF clusters were earlier also observed under the Frontier Fields Survey program in three HST-ACS/WFC (F435W, F606W, and F814W) and four HST-WFC3/IR (F105W, F125W, F140W, and F160W) filters, complementing previous observations and reaching a $5\sigma$ limiting magnitude of ${\sim}29$ AB (Lotz et al., 2017). We used these observations to measure the photometry and derive relevant stellar properties for each arc, and to estimate the microlens density in the vicinity of the arcs. For these purposes we used the Bayesian Analysis of Galaxies for Physical Inference and Parameter EStimation (bagpipes; Carnall et al., 2018) and spisea (Hosek et al., 2020) codes. More details are given in the section below.
Figure 1: Lensed arcs studied in our current work. In each panel, the name of the galaxy cluster, lensed arc system number (same as shown in Table LABEL:tab:arcs), and arc redshift are given in the upper-left part. The red curves show the critical curve at the arc redshift corresponding to the lens model (mentioned in the upper-right part of the panel) used in the current work for various estimates for each arc. The green rectangles mark apertures used to estimate the ICL density near each arc. The color images are made using HFF data with R=F105W+F125W+140W+160W; G=F606W+F814W+F105W; B=F435W+F606W.
3 Methodology for estimating the CCE rate
@@
3.1 Sample of lensed arcs
Lensed arcs used in the current work are listed in Table LABEL:tab:arcs, and the corresponding color stamps are shown in Figure 1. We studied a total of seventeen cluster caustic-crossing arcs in Flashlights. The twenty transients detected in the Flashlights program during the two visits are only found in seven of these arcs, whereas the remaining ten did not yield any transients. We refer readers to Kelly et al. (2022) for a view of the Flashlights difference imaging showing the transients in the various lensed arcs.
While this does not affect our analysis or results, we note that this is not a complete list of lensed arcs around cosmic noon in the HFF clusters. For example, there are some other lensed arcs that lie very close to one of the cluster galaxies such that they are heavily contaminated (complicating the analysis and likely also reducing the chances for observing CCEs in them), or pairs of lensed galaxy images that do not necessarily form arcs but can have certain parts with magnification factors $>10$, which may be sufficient to give rise to observable CCEs from very bright stars. Although their contribution is not expected to be major, a more exhaustive analysis should also consider such systems and is left for future work.
@@ 3.2 Microlens density
Before estimating the stellar population in the source galaxies, one important ingredient for calculating the rate of CCEs is the microlens density in the ICM close to these arcs. We placed multiple rectangular apertures near each lensed arc (see Figure 1) to photometrically measure the intracluster light (ICL) in the seven HFF filters. We used bagpipes to fit a spectral energy distribution (SED) to the photometry and estimate the ICL surface density near the lensed arcs. bagpipes uses simple stellar population synthesis (SSPS) models with the Kroupa IMF (Kroupa & Boily, 2002) along with an exponentially declining star-formation history (SFH) to do this. The measured stellar microlens density near each lensed arc is shown in column (5) of Table LABEL:tab:arcs, and we used the best-fit values for various estimations.